OpenStudy (chloelechloe):

Find the vertex of the graph of the function. f(x) = (x - 10)2 - 8

OpenStudy (3mar):

Could I help you?

OpenStudy (3mar):

If you still need help, let me know!

OpenStudy (nesha97):

the vertex is (10,-8)

OpenStudy (nesha97):

sorry if im late

OpenStudy (mathmale):

the function is actually y=(x-10)^2-8. Must use the operator " ^ " to denote exponentiation. Compare this to the vertex form of the equation of a parabola: y=(x-h)^2+k You can see that h=10, k=-8. From this you know that the vertex is at (10,-8).

OpenStudy (mathmale):

You need to know the standard equations of a parabola, to which you can compare any given equation of a parabola to determine the parameters (which, in this case, are h and k, the x- and y-coordinates of the vertex.